OpenStudy (anonymous):

evaluate the expression. (16/25)^3/2 3/2 is fraction @-------College Al…

OpenStudy (asnaseer):

\[a^{\frac{3}{2}}=(\sqrt{a})^3\]\[16=4^2\]\[25=5^2\]you should be able to solve this now...

OpenStudy (anonymous):

\[(\frac{16}{25})^\frac{3}{2}=\sqrt{(\frac{16}{25})^3}\]

OpenStudy (stacey):

The bottom of a fraction is the root you want and the other number is a normal exponent.